OpenStudy (anonymous):

SOMEONE PLS HELP ME !!! :((( 1) Write the rule for the nth term of the arithmetic sequence. 3, (-1) , (-5) , . . . 2) write a rule for the nth term of the arithmetic sequence with a_1=-7 and the common difference of 5/2 .

OpenStudy (babalooo):

An=a1+(n-1)d a1=3 d=-4 so it should be An = 3+(n-1)(-4). For number 2, just follow the formula to answer.